OpenStudy (anonymous):

Find an equation of a plane through the point (4, -4, 1) which is orthogonal to the line x=-4+5t, y=4-3t, z=-2+4t, in which the coefficient of x is 5.

OpenStudy (turingtest):

I'm not sure of the orthodox way to do this, but I think I got the solution...

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

the vector that points in the line given is (5,-3,4), agreed?

OpenStudy (anonymous):

yes i agree

OpenStudy (turingtest):

sorry wait I have to rethink :S

OpenStudy (anonymous):

its ok

OpenStudy (turingtest):

okay...since that vector is to be perpendicular to the plane, we can call it our normal vector n=(5,-3,4) now we need a vector in the plane. the position vector of the given point is Po=(4,-4,1) and we can call any other point in the plane P=(x,y,z) the vector P-Po will therefor be in the plane, which means that n(dot)(P-Po)=0

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

that translates to (5,-3,4)(dot)(x-4,y+4,z-1)=0 5(x-4)-3(y+4)+4(z-1)=0 5x-20-3y-12+4z-4=0 5x-3y+4z=36 I think I did that right, but it has been a while if it makes sense to you though it's probably good... I hope ;)

OpenStudy (anonymous):

yes it makes sense thank you

OpenStudy (anonymous):

Its correct thank you :)

OpenStudy (turingtest):

awesome :) very welcome, thanks for the practice!